OpenStudy (anonymous):

if the area formula for right angle triangles is base times height divided by 2 then if the base is 62 cm and the total area is 186 cm what is the height

OpenStudy (anonymous):

A = 186 = bh/2 = (62)(h)/2 h = (2)(186)/(62) = 6

OpenStudy (anonymous):

Checking: (62)(6)/2 = 186

OpenStudy (anonymous):

thank you for the answer

OpenStudy (anonymous):

You're welcome!